When hydrochloric acid is added to sodium carbonate, the reaction fizzes. Write a balanced chemical equation and predict how many grams of the product would remain after several days if 0.50g of sodium carbonate reacted completely.
1
Expert's answer
2017-03-08T08:41:06-0500
Balanced chemical equation: Na2CO3 + 2HCl = 2NaCl + H2O + CO2↑
ϑ(NaCl) = 2∙ϑ(Na2CO3)
m(NaCl) = M(NaCl)∙ϑ(NaCl) = 2∙M(NaCl)∙ϑ(Na2CO3) = 2∙M(NaCl)∙m(Na2CO3)/M(Na2CO3) = 2∙58.5∙0.5/106 = 0.552 (g) – mass of the product